Chair: Jay

Next meeting chair: Earl

Attended: Japser, Eloise, Lim, Mas, Jamie, Simin is here

Apologies: Prad

General Issues:

* Can't send the end exp. message - should be okay now * mk5cn timeout errors on yg - at the moment, it seems going on. Jamie will look into it. till then, keep an eye on that till now. * delay corrs generally being bad for Hb- nothing much we can do. needs power cyclying more and more and hope that it works.


Hobart 12m

Notes for wiki : after fs-aum command, it is not mentioned in wiki whether you need to select flexbuffhb or flexbuff-hb2, had to confer with Warren (flexbuff-hb2). It is also not mentioned which version of dbbc3hb to run, or whether or not a version of dbbchb must also be running. (Eloise) - flexbuff-2hb, dbbc3hb should be right choice.

No sampler “good/bad” readings - sampler summary script (multicast) not installed on pcfshb, and “AUM monitor” on ops8 not working either: - the multicast3*.py script is on pcfs-2hb. we can run it from there

R41019 Yarragadee 12m

A few mk5cn time-out errors which seem benign (Lim). 1156UT mk5cn time-out errors. Back on schedule from 1606UT (Jay). -most prob issue with the module. try to swap the module early on when you see error like this.

Si1281 Hobart 12m

Bad Luck. Most of Delay Corr went bad as soon as the Exp. started. No point of reconfig. the dbbc again because it will take ~30m and I doubt that it will get good even after that.

AOV064 Hobart 12m

2127UT Delay Corr A D E F gone bad (Lim) Issue is temperature in server room - AC under repair. Temporary ventilation in place but likely to have issues until AC restored. (JMc). First good data after restarting DBBC3 from 0506 (JMc) - one of a time mechanical issue

1313UT -restarted ddbc3 because sampler E was bad. The stats for other samplers worsened after restarting dbbc3. (Prad) - try power cycling it? we only need samplers A,B,C, D. so we can afford sampler E being bad

Katherine 12m

1655UT “Delay Corr” issue, experiment started late (Lim). Some nuisance errors concerning samplers being marginally below threshold. Issue seems to be transient (JMc) - Jamie will look into log monitor settings for this